The Marvosym F nt Package

Thomas Henlich (thenlich@rcs.urz.tu-dresden.de)

6th September 1998

This document describes how to use the Martin Vogels Symbole font in your L^{A}TEX documents. The font is a PS font, so you need a PS printer or ghostscript. The font provides the “original” Euro currency symbol, Times/Helvetica/Courier Euro symols, a CE symbol, astronomy symbols, zodiac signs and others.

1 Installation

In the following, replace TEXMF with your actual texmf directory, e. g. /usr/lib/teTeX/

texmf.

• Copy marvosym.pfb into a directory where dvips can find it, e. g. TEXMF/fonts/

type1/public/marvosym.

• Copy fmvr8x.tfm into a directory where TEX will find it, e. g. TEXMF/fonts/

tfm/public/marvosym.

• Add the line

fmvr8x␣Martin_Vogels_Symbole␣<marvosym.pfb
to your psfonts.map file, this should be TEXMF/dvips/misc/psfonts.map.

(Note: ␣ is a space, but _is an underline)

• Copy marvosym.sty to a TEX input directory.

• Optionally copy marvosym.afm into an afm-directory, e. g. TEXMF/fonts/afm/

public/marvosym.

2 Usage

In your document’s preamble, type \usepackage{marvosym}. To insert a symbol, use a macro as described in table ??. The symbol will appear in the currently selected size. To get

a bigger symbol, use a command like {\LARGE\Womanface}, which will produce .

   \Beam           \Bearing        \Loosebearing       \Fixedbearing      \Lefttorque      \Righttorque        \Force        \Lineload
\Corresponds       \Rightarrow     \Vectorarrowhigh       \Vectorarrow
\Rightscissors        \Kutline        \Leftscissors                                             \Cutright        \Cutline       \Cutleft
   \EUR                                \EURtm              \EURhv            \EURcr
  \Pickup           \Letter            \CEsign           \Mobilefone        \Industry        \Coffeecup       \Yingyang       \Telefon
\Clocklogo        \Checkedbox        \Crossedbox        \Pointinghand      \Writinghand       \Football          \fax           \FAX
\Faxmachine        \Snowflake            \Kross            \Lightning         \Circles           \Cross        \Celtcross        \Ankh
  \Heart            \Frowny            \Smiley             \Bicycle          \Stopsign
   \Sun              \Moon             \Mercury              \Venus             \Mars           \Jupiter         \Saturn        \Uranus
 \Neptune           \Pluto
\Zodiac{1}        \Zodiac{2}         \Zodiac{3}          \Zodiac{4}        \Zodiac{5}        \Zodiac{6}      \Zodiac{7}     \Zodiac{8}
  \Aries            \Taurus            \Gemini             \Cancer             \Leo            \Virgo          \Libra        \Scorpio
\Zodiac{9}        \Zodiac{10}        \Zodiac{11}         \Zodiac{12}
\Sagittarius       \Capricorn          \Aquarius             \Pisces
\Flatsteel        \Squarepipe         \Rectpipe             \Lsteel           \TTsteel         \Circpipe         \Tsteel
\Womanside         \Manside          \Womanfront           \Manfront            \Bat          \Womanface

Table 1: Marvosym macros

3 Authors

The font was designed by Martin Vogel (martin.vogel@fh-bochum.de). See http://www.

fh-bochum.de/fb1/vogel/marvosym.html.

Conversion to Type 1 font and the macros were done by Thomas Henlich (thenlich@rcs.urz.tudresden.de).

4 History

1998-07-20 Changed (*) to /* in /FontName. Thanks to Denis B. Roegel for telling me about this.

1998-06-21 Conversion to type 1 font now done with ttf2pt1 program. Font works now with dvips 5.78 and partial font downloading. Thanks to Uwe W. Gehring and Armin Geisse for cooperation. Added \Ankh macro. Renamed some macros.

1998-06-10 First version.